OpenStudy (anonymous):

If a tennis ball has a diameter of approximately 6.9 centimeters, what is its surface area

OpenStudy (anonymous):

do you know the formula for surface area of a sphere? (ball)

OpenStudy (anonymous):

I sure dont

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

Because I dont? thats why I want help ?

OpenStudy (anonymous):

kidding... here's the sufrace area formula: \[\large SA = 4 \times \pi \times r^2\]

OpenStudy (anonymous):

since your diameter is 6.9, take half of that to get your radius.... sorry gotta go... :(

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

Dpainc is right. Just use that formula.

OpenStudy (anonymous):

\[\text{Surface area} = 4*\pi*r^2\] And of course, if you're given the diameter instead, it's easy to find the radius from that.

OpenStudy (anonymous):

Yeahhh about that ... Im not to good in math so can you dumb it down a little more like go step by step with me please ? :)

OpenStudy (anonymous):

Step 1, figure out what the radius is. Usually it will be given to you. If they give you the diameter, the radius is always half of the diameter. Step 2, square the radius Step 3, multiply that by 4 step 5, multiply that by pi

OpenStudy (anonymous):

23.3 ?

OpenStudy (anonymous):

jk that isnt even an option on this test

OpenStudy (anonymous):

those are the choices they give me 149.57 square centimeters 1,032.04 square centimeters 299.14 square centimeters 598.28 square centimeters

OpenStudy (anonymous):

Is that a guess or did you try?

OpenStudy (anonymous):

I actually tried :/

OpenStudy (anonymous):

Okay, step by step then. What did you get as the radius?

OpenStudy (anonymous):

3.45

OpenStudy (anonymous):

Good. Show me the second step.

OpenStudy (anonymous):

149.57 thats the answer !

OpenStudy (anonymous):

Haha what did you do wrong the first time?

OpenStudy (anonymous):

I took the square root instead of just squaring it

OpenStudy (anonymous):

Oooooh =D

OpenStudy (anonymous):

Yes but I still need your help so dont leave ! lol

OpenStudy (anonymous):

If the smaller cylinder has a lateral area of approximately 113 square meters, what is the approximate lateral area of the larger cylinder The bigger cylinder is 21m an the smaller one is 9m

OpenStudy (anonymous):

Okay, well this problem has a few steps, but we're really only going to use 2 facts 1) Lateral surface area = pi*diameter*height 2) Similar shapes have corresponding parts that are proportional. Here's the picture we have